1. order: phenobarbital sodium 400mg po in 2 divided dose daily. the recommended dosage for this anticonvulsant is 1-3mg/kg/day. is the prescribed dose is safe for the patient who weighs 203 pounds.

2. Order: proventil 2mg PO T.I.D. the recommended dosage range Is 0.1-0.2mg/kg tid max 4mg/dose. the child weighs 32 pounds. the label read 2mg/5ml. if the dose is safe, how many teaspoon of this bronchodilator will the child receives.

need step by step process to get the answers.
Answer #1

1- 203 pounds waight= 92 kg

Maximum dose give = 92× 3= 276mg in a day

So 400 mg is not safe for this patient

2- child waight 32pounds=14kg

1.4-2.8 mg tid for this child use , so dose is safe because not exceed greater than 4 mg per dose

1.4 - 2.8 average 2 mg given

2mg in 5 ml

So tid - means 3 spoon given